OpenStudy (anonymous):

For the values a = 2.9 which is a leg of a right triangle, and value c = 4.2, which is the hypotenuse, find the length of the other leg, b, to the nearest tenth

OpenStudy (anonymous):

Pythagorean theorem a^2+b^2=c^2 so: 2.9^2+b^2=4.2^2 solve for b

OpenStudy (anonymous):

1.3